40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 27 15:18:50 2011, kiwamu, Summary, General, schedule 
    Reply  Wed Jul 27 15:42:41 2011, kiwamu, Summary, General, mitues of 40m meeting : Task list 
       Reply  Sun Jul 31 22:48:35 2011, kiwamu, Summary, General, preparation of the vent : status update 
          Reply  Mon Aug 1 20:21:05 2011, kiwamu, Update, Green Locking, Y arm green beam axis : aligned 
             Reply  Tue Aug 2 02:20:09 2011, kiwamu, Update, Green Locking, Y arm green beam axis : done 
Message ID: 5084     Entry time: Mon Aug 1 20:21:05 2011     In reply to: 5078     Reply to this: 5088
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: Y arm green beam axis : aligned 

The beam axis of the Y green light has been aligned.

Now I can see TEM00 mode is flashing on the ETMY camera.

 

-- (What I will do tonight)

The next step is to refine some electronics in the PDH loops to get the green light locked to the Y arm cavity.

If the beam isn't locked, I guess the in-vac-work will be so difficult because of the low intensity of the green light.

According to a brief check on the circuits, a low pass filter after the demodulation mixer is in a sad situation.

It doesn't pass any signals and in fact it behaves more like an absorber.

On the other hand, the modulation system looks fine to me because I was able to see the 270 kHz sideband converted into AM due to the fringing.

Quote from #5078

  (not yet) Alignment of the Y green beam (#5066)

ELOG V3.1.3-